Answer the tollowing questions. Table 6-4 or Table 6-5. (Use appropriate factor(s) from the tables provided. Round the PV factors to 4 decimals.) Required: a. Spencer Co.s common stock is expected to have a dividend of $6 per share for cach of the next twelve years, and it is estimated that the market value per share will be S133 at the end of twelve years. If an investor requires a return on investment of 4%, what is the maximum price the investor would be willing to pay for a share of Spencer Co. common stock today? Do not round intermediate calculations. Round your answer to 2 decimal places b Mario bought a bond with a face amount of $1,000 a stated interest rate of 6% and a maturity date twenty years in the future or $984. The bond pays interest on an anual basis. Three years have gone by and the market interest rate is now 12% What is the market value of the bond today? (Do not round intermediate calculations. Round your answer to 2 decimal places.) c. Alexis purchased a U.S. Series EE savings bond for $100, and six years later recoived $141.84 when the bond was redeemed. What average annual retum on investment did Alexis earn over the sox years?TABLE 6.4 FACTORS FOR CALCULATING THE PRESENT VALUE OF $1 Discount Rate No. of Periods 2% 4°10 6% 8% 10% 12% 1 4% 1 6% 18% 20% 0.980 0.9615 0.9434 0.9259 0.9091 0.8929 0.8772 0.8621 0.8475 0.8333 0.961 0.9246 0.8900 0.8573 0.8264 0.7972 0.7695 0.7432 0.7182 0.6944 0.942 0.8890 0.8396 0.7938 0.7513 0.7118 0.6750 0.6407 0.6086 0.5787 0.924 0.8548 0.7921 0.7350 0.6830 0.6355 0.5921 0.906 0.8219 0.7473 0.6806 0.6209 0.5674 0.5194 0.4761 0.888 0.7903 0.7050 0.6302 0.5645 0.5066 0.4556 0.4104 0.3704 0.3349 0.871 0.7599 0.6651 0.5835 0.5132 0.4523 0.3996 0.3538 0.3139 0.2791 0.853 0.7307 0.6274 0.5403 0.4665 0.4039 0.3506 0.3050 0.2660 0.2326 0.837 0.7026 0.5919 0.5002 0.4241 0.3606 0.3075 0.2630 0.2255 0.1938 0.820 0.6756 0.5584 0.4632 0.3855 0.3220 0.2697 0.2267 0.1911 0.1615 0.804 0.6496 0.5268 0.4289 0.3505 0.2875 0.2366 0.1954 0.1619 0.1346 0.788 0.6246 0.4970 0.3971 0.3186 0.2567 0.2076 0.1685 0.1372 0.1122 0.773 0.6006 0.4688 0.3677 0.2897 0.2292 0.1821 0.1452 0.1163 0.0935 0.758 0.5775 0.4423 0.3405 0.2633 0.2046 0.1597 0.1252 0.0985 0.0779 0.743 0.5553 0.4173 0.3152 0.2394 0.1827 0.1401 0.1079 0.0835 0.0649 0.728 0.5339 0.3936 0.2919 0.2176 0.1631 0.1229 0.0930 0.0708 0.0541 0.714 0.5134 0.3714 0.2703 0.1978 0.1456 0.1078 0.0802 0.0600 0.0451 0.700 0.4936 0.3503 0.2502 0.1799 0.1300 0.0946 0.0691 0.0508 0.0376 0.686 0.4746 0.3305 0.2317 0.1635 0.1161 0.0829 0.0596 0.0431 0.0313 0.673 0.4564 0.3118 0.2145 0.1486 0.1037 0.0728 0.0514 0.0365 0.0261 0.660 0.4388 0.2942 0.1987 0.1351 0.0926 0.0638 0.0443 0.0309 0.0217 0.647 0.4220 0.2775 0.1839 0.1228 0.0826 0.0560 0.0382 0.0262 0.018 0.634 0.4057 0.2618 0.1703 0.1117 0.0738 0.0491 0.0329 0.0222 0.015 0.622 0.3901 0.2470 0.1577 0.1015 0.0659 0.0431 0.0284 0.0188 0.0126 0.610 0.3751 0.2330 0.1460 0.0923 0.0588 0.0378 0.0245 0.0160 0.0105 0.552 0.3083 0.1741 0.0994 0.0573 0.0334 0.0196 0.0116 0.0070 0.0042 0.500 0.2534 0.1301 0.0676 0.0356 0.0189 0.0102 0.0055 0.0030 0.0017 0.453 0.2083 0.0972 0.0460 0.0221 0.0107 0.0053 0.0026 0.0013 0.0007 0.410 0.1712 0.0727 0.0313 0.0137 0.0061 0.0027 0.0013 0.0006 0.0003 0.372 0.1407 0.0543 0.0213 0.0085 0.0035 0.0014 0.0006 0.0003 0.0001 3 4 0.5523 0.5158 0.4823 0.4371 0.4019 6 7 8 10 12 13 15 17 18 20 23 24 25 30 35 40 45 50

TABLE 6.5 FACTORS FOR CALCULATING THE PRESENT VALUE OF AN ANNUITY OF $1 Discount Rate No. of Periods 2% 4% BY。 8% 10% 12% 14% 1 6% 18% 20% 0.980 0.9615 0.9434 0.9259 0.9091 0.8929 0.8772 0.8621 0.8475 0.8333 1.942 1.8861 1.8334 1.7833 1.7355 1.6901 16467 6052 5656 5278 2.884 2.7751 2.6730 2.5771 2.4869 2.4018 2.3216 2.2459 2.1743 2.1065 3.808 3.6299 3.4651 3.3121 3.1699 3.0373 2.9137 2.7982 2.6901 2.5887 3.2743 3.1272 2.9906 3 4.713 4.4518 4.2124 3.9927 3.7908 3.6048 3.4331 6 5.601 5.2421 4.9173 4.6229 4.3553 4.1114 3.8887 3.68473.4976 3.3255 6.472 6.0021 5.5824 5.2064 4.8684 4.5638 4.2883 4.0386 3.8115 3.6046 7.325 6.7327 6.2098 5.7466 5.3349 4.9676 4.6389 4.3436 4.0776 3.8372 8.162 7.4353 6.8017 6.2469 5.7590 5.3282 4.9464 4.6065 4.3030 4.0310 8.983 8.1109 7.3601 6.7101 6.1446 5.6502 5.2161 4.8332 4.4941 4.1925 10 9.787 8.7605 7.8869 7.1390 6.4951 5.9377 5.4527 5.0286 4.6560 4.3271 10.575 9.3851 8.3838 7.5361 6.8137 6.1944 5.6603 5.1971 4.7932 4.4392 11.348 9.9856 8.8527 7.9038 7.1034 6.4235 5.8424 5.3423 4.9095 4.5327 12.106 10.5631 9.2950 8.2442 7.3667 6.6282 6.0021 5.4675 5.0081 4.6106 12.849 11.1184 9.7122 8.5595 7.6061 6.8109 6.1422 5.5755 5.0916 4.6755 15 13.578 11.6523 10.1059 8.8514 7.8237 6.9740 6.2651 5.6685 5.1624 4.7296 14.292 12.1657 10.4773 9.1216 8.0216 7.1196 6.3729 5.7487 5.2223 4.7746 14.992 12.6593 10.8276 9.3719 8.2014 7.2497 6.4674 5.8178 5.2732 4.8122 15.678 13.1339 .1581 9.6036 8.3649 7.3658 6.5504 5.8775 5.3162 4.8435 16.351 13.5903 11.4699 9.8181 8.5136 7.4694 6.6231 5.9288 5.3527 4.8696 20 23 24 25 17.011 14.0292 11.7641 10.0168 8.6487 7.5620 6.6870 5.9731 5.3837 4.8913 17.658 14.4511 12.0416 10.2007 8.7715 7.6446 6.7429 6.0113 5.4099 4.9094 18.292 14.8568 12.3034 10.3711 8.8832 7.7184 6.7921 6.0442 5.4321 4.9245 18.914 15.2470 12.5504 10.5288 8.9847 7.7843 6.8351 6.0726 5.4509 4.9371 19.523 15.6221 12.7834 10.6748 9.0770 7.8431 6.8729 6.0971 5.4669 4.9476 30 35 40 45 50 22.396 17.2920 13.7648 11.2578 9.4269 8.0552 7.0027 6.1772 5.5168 4.9789 24.999 18.6646 14.4982 11.6546 9.6442 8.1755 7.0700 6.2153 5.5386 4.9915 27.355 19.7928 15.0463 11.9246 9.7791 8.2438 7.1050 6.2335 5.5482 4.9966 29.490 20.7200 15.4558 12.1084 9.8628 8.2825 7.1232 6.2421 5.5523 4.9986 31.424 21.4822 15.7619 12.2335 9.9148 8.3045 7.1327 6.2463 5.5541 4.9995
Answer #1

a. Present value of dividend of $6 receivable for next twelve years @4% = $6*9.3851 = $56.31

Present value of $133 at the end of twelve years = $133*0.6246 = $83.07

So, the maximum price the investor will pay for the share = $56.31 + $83.07 = $139.38

b. Bond is of 20 years and 3 years have already been passe, so the remaining life of the bond is 17 years.

Present value of interest of $60 receivable for next seventeen years @12% = $60*7.1196 = $427.18

Present value of $1000 at the end of seventeen years = $1000*0.1456 = $145.60

So, the market value of the bond = $427.18 + $145.60 = $572.78

c. Alexis received $141.84 after 6 years by investing $100.00 in savings bond. The formula for future value = PV*(1+r)^n

Here 141.84 = 100*(1+r)^n

So (1+r)^n = 1.4184

So for r=6%, (1+r)^n = 1.4184

So the average annual return on investment here is 6%.